how much water must be added to 500g of 30 acid solution in order to produce a 20% acid solution?
Answer by stanbon(75887)

how much water must be added to 500g of 30 acid solution in order to produce a 20% acid solution?
----
Equation:
water + water = water
0.70*500 + x = 0.80(500+x)
------
70*500 + 100x = 80*500 + 80x
20x = 10*500
x = 250 g (amt. of water needed)
====================
Cheers,
Stan H.
------------------------